Probability distribution of discrete random variable x is given as:
x -2 -1 1 2
p(X=x) .3 .3 .1 .3
Find P(x is an even number)

Answer by stanbon(75887)

Probability distribution of discrete random variable x is given as:
x -2 -1 1 2
p(X=x) .3 .3 .1 .3
Find P(x is an even number) = P(x = -2 or x = 2) = 0.3+0.3 = 0.6
===================
Cheers,
Stan H.
=====================